Hydrogen produced from a hydrolysis reaction was collected over water. The total volume of h2 (g) was 93.26 at the temperature of 25 degree C. The barometric pressure was 742 mmHg and the vapor pressure of water at 25 degrees C was 24.0 mmHg. Calculate the moles of hydrogen gas produced by the reaction


1
Expert's answer
2021-02-03T07:15:48-0500

Volume(V) = 93.26 mL = 93.26/1000= 0.09326 L

T = 25 °C = 25 + 273 = 298 K

P = 742 - 24.0 = 718 mm Hg = 718/760 = 0.945 atm

Using ideal gas equation,

PV = nRT

Where, P = pressure

V = Volume

n = mole

R = Gas constant

T = Temperature

putting the values

0.945 × 0.09326 = n × 0.0821 × 298

n = 0.00360 mol

So, moles of Hydrogen gas produced = [0.00360] mol